Find the probability that the sum of the values of two dice in a throw of two fair dice is 8, given that the one die is 5.
Answer by drk(1908)

Given one die is 5 we could have (5,3) or (3,5). So,
P (sum = 8, given one is 5) = 2/36 = 1/18.